28. Balfour corporation has a current stock price of $30. Next year's dividend is projected to be $3.00. The payout ratio is 30% and projected ROE is 10%. the cost of equity is:

  1. 16%

  2. 17%

  3. 18%

  4. None of the above

31. The present value of future cash flow per share is $50 for rose corporation. the current earnings-per-share is $5.00. the implied price-earnings ratio is:

  1. 5

  2. 10

  3. 15

  4. None of the above
Answer #1

28)

Dividend 3
Present stock price 30
Payout ratio 0.3
Retention Ratio (RR) 0.7
ROE 10%
Growth Rate 7% (ROE*RR )

Ke = (d/s)+g = 17%

31) Implied earnings ratio = 5/5*100 = 10%